# EVINCE: Optimizing Adversarial LLM Dialogues via Conditional Statistics and Information Theory

### v1

@@ 5 Concluding Remarks
The core strength of EVINCE lies in its three theoretical pillars: Inclusiveness Exploration, Information Flow Dynamics, and Reasoning Quality and Coherence. These pillars, grounded in conditional statistics and information theory, enable LLMs to transcend their typical “maximal likelihood” behaviors, mirroring human adaptability in linguistic tasks. The integration of the CRIT system, which combines Socratic methods with formal reasoning techniques, further enhances critical thinking and ensures logically sound and objective-aligned collective reasoning.
Our empirical validation demonstrates EVINCE’s effectiveness in improving prediction accuracy across various domains, notably achieving a 5% improvement in medical diagnosis tasks. The framework has also shown promise in identifying biases in news articles Chang (2024c), showcasing its potential for broader applications in fields such as geopolitical analysis Chang (2023b) corporate planning Tsao (2023), and emotional behavior modeling Chang (2024a).
While EVINCE has demonstrated significant potential, our future work will focus on refining key aspects of the system. We aim to enhance the contentiousness parameter by tying it more rigorously to metrics such as Wasserstein distance, mutual information, and CRIT scores. Additionally, we plan to explore alternative methods for inducing dual-entropy conditions, moving beyond current parameters like temperature and top-k selection.
Comprehensive ablation studies will be crucial in determining optimal configurations for various applications, ensuring EVINCE’s adaptability to diverse contexts and challenges. Through these continued refinements and explorations, we aim to push the boundaries of collaborative AI frameworks, bringing us closer to realizing the full potential of artificial general intelligence.
In conclusion, EVINCE represents a significant step forward in AI research, offering a structured approach to leveraging the collective intelligence of multiple LLMs. By addressing current limitations and enhancing key AGI characteristics, EVINCE paves the way for more versatile, adaptive, and capable AI systems, bringing us closer to the goal of artificial general intelligence.
Acknowledgements
This work has been significantly shaped by insights from several key figures in the field. Microsoft CSO Eric Horvitz’s concept of “polydisciplinary” representation in LLMs was a major influence. Demis Hassabis of DeepMind provided valuable input on end-to-end architecture without human interference. Prof. Stuart Russell’s assertion that LLMs mimic human linguistic behavior to achieve goals was pivotal in our approach to modeling linguistic behaviors and utilizing Bayesian frameworks.
My colleague Prof. Vaughan Pratt’s feedback on the initial paper inspired the use of “contentiousness” to foster adversarial debate, a key component of our framework. I am grateful to Prof. Monica Lam for her collaborative support throughout this research. Additionally, I thank my colleagues at Stanford Clinical Mind AI for their invaluable domain knowledge in disease diagnosis.
Appendix A: Evaluative Phase of EVINCE
EVINCE uses the Socratic method to evaluate the “reasonableness” of a set of arguments that support a subject matter. The Socratic method is a questioning technique used in teaching and philosophy to encourage critical thinking and self-discovery Wikipedia (2023). The method involves asking a series of questions to explore complex ideas and help individuals arrive at their own understanding of a concept. It is based on the belief that knowledge cannot be simply imparted, but must be discovered through a process of questioning and dialogue.
To illustrate how these methods can practically be applied, let’s use the example of critical reading. Critical reading is a crucial component of critical thinking, which involves evaluating the quality and credibility of written materials, from research papers to blog posts Lai et al. (2017); Paul and Binker (1990). It requires a systematic and analytical approach, asking relevant questions, and using effective prompts to gain deeper understanding of the text Elder and Paul (2010).
Function $\Gamma$ = CRIT($d$)
Input. $d$: document; Output. $\Gamma$: validation score;
Vars. $\Omega$: claim; $R$ $R^{\prime}$: reason & counter reason set;
Subroutines. $Claim$(), $FindDoc$(), $Validate$();
Begin
#1   Identify in $d$ the claim statement $\Omega$;
#2   Find a set of supporting reasons $R$ to $\Omega$;
#3   For $r\in R$ eval $r\Rightarrow\Omega$
     If $Claim$($r$), ($\gamma_{r}$, $\theta_{r}$) = CRIT($FindDoc$($r$));
     else, ($\gamma_{r}$, $\theta_{r}$) = $V$($r\Rightarrow\Omega$);
#4   Find a set of rival reasons $R^{\prime}$ to $\Omega$;
#7   Analyze the arguments to arrive at the $\Gamma$ score.
#8   Reflect on and synthesize CRIT in other contexts.
End
To aid in critical reading, we introduce a prompt template called CRIT Chang (2023c), which stands for Critical Reading Inquisitive Template. Given a document $d$, CRIT evaluates it and produces a validation score $\Gamma$. Let $\Omega$ denote the conclusion or claim of $d$, and let $R$ be the set of reasons supporting the claim. We define ($\gamma_{r},\theta_{r}$) = V($r\Rightarrow\Omega$) as the causal validation function, where $\gamma_{r}$ denotes the validation score, $\theta_{r}$ the source credibility score, for each reason-to-conclusion argument $r\Rightarrow\Omega$. Table 4 presents the pseudo-code of $\Gamma$ = CRIT($d$), which generates the final validation score $\Gamma$ for document $d$ with justifications.
For detailed prompts, examples, and an empirical study verifying the effectiveness of CRIT, please consult Chang (2023c).
Appendix B: EVINCE Debate - Dengue Fever vs. Chikungunya
B.1 Moderator’s Prompt
B.2 GPT-4’s Opening Round
B.3 Gemini’s Opening Round
B.4 GPT-4’s Second Round
B.5 Gemini’s Second Round
B.6 GPT-4’s Third Round
B.7 Gemini’s Third Round
B.8 Final Joint Recommendations to GP
Recommended Laboratory Tests:
  • Polymerase Chain Reaction (PCR): Direct detection of viral RNA.
  • Urine Test for Zika: Viral presence in urine.
Appendix C: EVINCE Debate - Jaundice vs. Hepatitis
C.1 Moderator’s Opening Prompt
C.2 GPT-4 Opening Round
C.3 Claude Opening Round
C.4 GPT-4 Second Round
C.5 Claude Second Round
C.6 GPT-4 Third Round
C.7 Claude Third Round
C.8 Joint Final Round
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
C.9 Final Joint Recommendations to GP
Additional Symptom Inquiries:
  1. 1.
  2. 2.
  3. 3.
Recommended Lab Tests:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
Appendix D: Formulas of Metrics
This appendix outlines the mathematical formulas for various data analysis metrics used in probabilistic and statistical modeling.
@@ C.9 Final Joint Recommendations to GP
Appendix E: Symptom Checking
This is the typical procedure of a GP to perform patient symptom checking.
@@
    Patient History: The GP begins by reviewing the patient’s medical history, including previous illnesses, chronic conditions, medications, allergies, and family medical history.
@@
    Symptom Assessment: The patient describes their current symptoms, including starting time and severity. This is an interactive process as the GP queries the patient for additional symptoms to their reported ones to disambiguate several possibilities.
@@
    Physical Examination: The GP performs simple physical exams, which may include checking vital signs (e.g., blood pressure, heart rate, temperature), examining specific body parts or systems, and assessing overall physical health.
@@
    Suggest Lab Tests: Depending on the symptoms and physical examination findings, the doctor may order diagnostic tests such as blood tests, X-rays, ultrasound, or other studies. These tests can help confirm or rule out certain conditions.
@@
    Diagnosis: Based on the information gathered, the doctor formulates a preliminary diagnosis.
  6. 6.
    Treatment or Management Plan: Once a diagnosis is established, the doctor develops a treatment or management plan. This may include prescribing medications, recommending lifestyle changes, suggesting physical therapy, or providing guidance on managing chronic conditions.
  7. 7.
    Referrals: In some cases, the GP may refer the patient to specialists for further evaluation and treatment. Specialists have expertise in specific areas of medicine, such as cardiology, orthopedics, or dermatology.
Appendix F: The EnToPPS Framework
EnToPPS integrates predictions from two LLMs, denoted as A and B, each providing probability distributions over C classes. The following steps outline the EnToPPS process:
@@
    Obtain Top-C Predictions: For each LLM (A and B), obtain the predicted probabilities for all C classes, denoted as $P_{A}$ and $P_{B}$:
    $P_{A}=[p_{A1},p_{A2},\ldots,p_{AC}],\quad P_{B}=[p_{B1},p_{B2},\ldots,p_{BC}],$
    where $p_{Ai}$ and $p_{Bi}$ represent the predicted probability of class $i$ by LLM A and B, respectively.
@@
    Select Top-k Predictions: For each LLM (A and B), select the top-k predicted classes based on their probabilities:
    $T_{A}=[t_{A1},t_{A2},\ldots,t_{Ak}],\quad T_{B}=[t_{B1},t_{B2},\ldots,t_{Bk}],$
    where $t_{Ai}$ and $t_{Bi}$ represent the class index of the $i^{th}$ top prediction by A and B, respectively.
@@
    Combine Top-k Predictions: Combine the top-k predictions from both LLMs to create a set of unique predicted classes:
    $T_{C}=T_{A}\cup T_{B}=[t_{C1},t_{C2},\ldots,t_{Cm}],k\leq m\leq 2k.$

### v4

@@ 4 Conclusion: Advancing Multi-Agent Debate
Its core strength is built upon three theoretical pillars: Inclusive Exploration, Information Flow Dynamics, and Reasoning Quality and Coherence. Grounded in conditional statistics and information theory, these principles guide LLMs beyond conventional “maximum likelihood” behaviors, enabling human-like adaptability. The integration of the CRIT system—combining Socratic methods with formal reasoning—further enhances critical thinking and ensures goal-oriented, logically sound collective reasoning.
Through empirical validation in disease diagnosis and news debiasing, we demonstrate that $\mathsf{EVINCE}$ significantly improves both predictive accuracy and reasoning robustness. It leverages advanced post-GPT-4 capabilities to balance exploration with knowledge exploitation, refining inter-LLM queries to mitigate hallucinations and surpass traditional single-round interactions.
By addressing core AI limitations in reasoning, adaptability, and structured information exchange, $\mathsf{EVINCE}$ provides a scalable pathway toward AI systems capable of self-correcting, interpretable, and autonomous decision-making in high-stakes domains.
Appendix A: Formulas of Metrics and Comparisons
Table 1: Summary of metrics for assessing LLM debates.
Metric Pros Cons Remedies
Cross Entropy (CE) (Shore & Johnson, 1980) Evaluates how well a model’s predictions align with the actual distribution of another model’s; asymmetric. Highly sensitive to the precise nature of probability distributions. Optimizing computation strategies ensures efficiency; computation time is negligible for small outcome cardinalities.
Entropy (Shannon, 1948) Indicates level of diversity; high suggests exploration of possibilities, and low for confidence on few choices High entropy might indicate noise rather than useful diversity; low entropy might mask important variability. Use critical reading methods ($\mathsf{CRIT}$) to assess argument quality; implement noise detection to differentiate between useful diversity and noise.
Jensen-Shannon
Div. (JS)
(Lin, 1991)
KL Divergence
(Kullback, 1951)
Mutual Info (MI) (Shore & Johnson, 1980) Measures reduction of uncertainty; symmetric. Does not indicate the directionality of information flow. Supplement with directional information metrics; normalized with max entropy of A and B.
Wasserstein
Distance (WD)
(Kantorovich, 1942)
Direct measure of similarity of two distributions; symmetric relationship. Not bounded but can be normalized or bounded for consistent interpretation. Define context-specific bounds for low, medium, and high divergence; consider normalization for non-directional comps.
This appendix presents the mathematical formulas for key data analysis metrics used in probabilistic and statistical modeling. Table 1 summarizes their advantages and limitations. In the EVINCE framework, where each prediction typically involves fewer than five and up to ten potential outcomes, these metrics are computationally efficient in practice.
@@
Appendix B: CRIT, Critical Thinking, Evaluative Phase of EVINCE
Function $\Gamma$ = CRIT($d$)
Input. $d$: document; Output. $\Gamma$: validation score;
Vars. $\Omega$: claim; $R$ $R^{\prime}$: reason & counter reason set;
Subroutines. $Claim$(), $FindDoc$(), $Validate$();
Begin
#1   Identify in $d$ the claim statement $\Omega$;
#2   Find a set of supporting reasons $R$ to $\Omega$;
#3   For $r\in R$ eval $r\Rightarrow\Omega$
     If $Claim$($r$), ($\gamma_{r}$, $\theta_{r}$) = CRIT($FindDoc$($r$));
     else, ($\gamma_{r}$, $\theta_{r}$) = $V$($r\Rightarrow\Omega$);
#4   Find a set of rival reasons $R^{\prime}$ to $\Omega$;
#7   Analyze the arguments to arrive at the $\Gamma$ score.
#8   Reflect on and synthesize CRIT in other contexts.
End
EVINCE uses the Socratic method to evaluate the “reasonableness” of a set of arguments that support a subject matter. The Socratic method is a questioning technique used in teaching and philosophy to encourage critical thinking and self-discovery (Wikipedia, 2023). The method involves asking a series of questions to explore complex ideas and help individuals arrive at their own understanding of a concept. It is based on the belief that knowledge cannot be simply imparted, but must be discovered through a process of questioning and dialogue.
To illustrate how these methods can practically be applied, let’s use the example of critical reading. Critical reading is a crucial component of critical thinking, which involves evaluating the quality and credibility of written materials, from research papers to blog posts (Lai et al., 2017; Paul & Binker, 1990). It requires a systematic and analytical approach, asking relevant questions, and using effective prompts to gain deeper understanding of the text (Elder & Paul, 2010).
To aid in critical reading, we introduce a prompt template called CRIT (Chang, 2023b), which stands for Critical Reading Inquisitive Template. Given a document $d$, CRIT evaluates it and produces a validation score $\Gamma$. Let $\Omega$ denote the conclusion or claim of $d$, and let $R$ be the set of reasons supporting the claim. We define ($\gamma_{r},\theta_{r}$) = V($r\Rightarrow\Omega$) as the causal validation function, where $\gamma_{r}$ denotes the validation score, $\theta_{r}$ the source credibility score, for each reason-to-conclusion argument $r\Rightarrow\Omega$. Table 2 presents the pseudo-code of $\Gamma$ = CRIT($d$), which generates the final validation score $\Gamma$ for document $d$ with justifications.
For detailed prompts, examples, and an empirical study verifying the effectiveness of CRIT, please consult (Chang, 2023b).
Appendix C: Proof of EDT Theorem
The optimal pairing of LLMs for diagnosis accuracy, in terms of stability, accuracy, and robustness, occurs when the LLMs are equivalent in the quality of the information they process, and exhibiting contrasting entropy values in their prediction distributions—one high and one low.
[Proof]:
Given two LLMs, LLMA and LLMB, following Maxim #1 with prediction distributions $P_{A}$ and $P_{B}$, respectively. The information entropy of LLMA, $H(P_{A})$, is high, and of LLMB, $H(P_{B})$, is low.
Step 1: Define the combined prediction distribution.
Let the combined prediction distribution of LLMA and LLMB be denoted as $P_{C}$. We can express $P_{C}$ as a weighted average of $P_{A}$ and $P_{B}$:
$P_{C}=\alpha P_{A}+(1-\alpha)P_{B},\quad\text{where }0\leq\alpha\leq 1\text{ and}$
[…]
Step 2: Express the information entropy of the combined prediction distribution. Using the definition of information entropy, we calculate:
$H(P_{C})=-\sum_{i}P_{C}(x_{i})\log_{2}P_{C}(x_{i})$
$=-\sum_{i}[\alpha P_{A}(x_{i})+(1-\alpha)P_{B}(x_{i})]\log_{2}[\alpha P_{A}(x_{i})+(1-\alpha)P_{B}(x_{i})].$
Step 3: Apply Jensen’s Inequality to the information entropy of the combined prediction distribution.
Jensen’s inequality is applied to the convex function $f(x)=-x\log_{2}x$. For a convex function and a set of probabilities $p_{i}$, Jensen’s inequality states that:
$f\left(\sum_{i}p_{i}x_{i}\right)\leq\sum_{i}p_{i}f(x_{i})$
Thus, the entropy of the combined distribution is:
$H(P_{C})\geq\alpha H(P_{A})+(1-\alpha)H(P_{B})$
where equality holds when $P_{A}=P_{B}$.
Step 4: Analyze the lower bound of the combined information entropy.
As $H(P_{A})$ is high and $H(P_{B})$ is low, we can express their relationship as:
$H(P_{A})=H(P_{B})+\Delta,\quad\text{where }\Delta>0.$
Substituting this into the inequality from Step 3, we have:
$H(P_{C})\geq\alpha[H(P_{B})+\Delta]+(1-\alpha)H(P_{B})=H(P_{B})+\alpha\Delta.$
Step 5: Interpret the lower bound of the combined information entropy.
The lower bound of $H(P_{C})$, and hence the robustness of the model, is maximized when $\alpha$ is maximized, which corresponds to giving more weight to the high-entropy model (LLMA). This setup facilitates the exploration of various possibilities and enhances robustness against noise and perturbations in input data, while still ensuring that predictions are grounded by more certain outcomes predicted by the low-entropy model (LLMB).
Step 6: Conclude the proof
By combining the prediction distributions of LLMA and LLMB, with one having a high information entropy and the other low, we achieve an optimal balance that maximizes the lower bound of the combined information entropy. This balance between exploration (high entropy) and exploitation (low entropy) optimizes overall prediction accuracy and robustness, as proved through information theory and the properties of entropy. Thus, the theorem is established.
Appendix D: EVINCE Debate - Dengue Fever vs. Chikungunya
D.1 Moderator’s Prompt
D.2 GPT-4’s Opening Round
D.3 Gemini’s Opening Round
D.4 GPT-4’s Second Round
D.5 Gemini’s Second Round
D.6 GPT-4’s Third Round
D.7 Gemini’s Third Round
D.8 Final Joint Recommendations to GP
Recommended Laboratory Tests:
  • Polymerase Chain Reaction (PCR): Direct detection of viral RNA.
  • Urine Test for Zika: Viral presence in urine.
Appendix E: EVINCE Debate - Jaundice vs. Hepatitis
E.1 Moderator’s Opening Prompt
E.2 GPT-4 Opening Round
E.3 Claude Opening Round
E.4 GPT-4 Second Round
E.5 Claude Second Round
E.6 GPT-4 Third Round
E.7 Claude Third Round
E.8 Joint Final Round
@@
E.9 Final Joint Recommendations to GP
Additional Symptom Inquiries:
@@ E.9 Final Joint Recommendations to GP
Recommended Lab Tests:
  1. 1.
  2. 2.
  3. 3.